On January 1, 20X5, Pirate Company acquired all of the outstanding stock of Ship Ic., a Norwegian company, at a cost of $167,Requlred: a. Prepare a schedule translating the trial balance from Norwegian kroner into U.S. dollars. Assume the krone is thb. Assume that Pirate uses the fully adjusted equity method. Record all journal entries that relate to its investment in thed. Compute Pirates total consolidated stockholders equity at December 31, 20X5. Consolidated stockholders equity
Answer #1

Part A

Items

Norweign Kroner

Exchange rate

U.S. Dollar

Cash

163000

0.21

34230

Accounts receivable (net)

222000

0.21

46620

Inventory

286000

0.21

60060

Property, plant and equipment

630000

0.21

132300

Cost of goods sold

411000

0.20

82200

Operating expenses

120000

0.20

24000

Depreciation expenses

67000

0.20

13400

Dividend paid

41000

0.19

7790

Total

1940000

400600

Accumulated depreciation[P1]

161000

0.21

33810

Accounts payable

97000

0.21

20370

Notes payable

191000

0.21

40110

Common stock

440000

0.18

79200

Retained earnings

260000

0.18

46800

Sales

791000

0.20

158200

Total

1940000

378490

Accumulated other comprehensive income – translation adjustment (Credit)

22110

Total credits

1940000

400600

Part B

No.

Date

General journal

Debit

Credit

1

January 1

Investment in Ship company

167400

Cash

167400

(to record purchase of Ship company)

2

July 1

Cash (41000*0.19)

7790

Investment in Ship company

7790

(to record dividend received from subsidiary)

3

December 31

Investment in Ship company (791000-411000-120000-67000)*0.20

38600

Income from Subsidiary

38600

(to record equity in net income of foreign subsidiary)

4

December 31

Investment in Ship company

22110

Other comprehensive income – Translation Adjustment

22110

(to record parent’s share of translation adjustment from translation of subsidiary’s accounts

5

December 31

Income from Subsidiary

7200

Investment in Ship company (2000+5200)

7200

(to record amortization)

6

December 31

Investment in Ship company (2900+3640)

6540

Other comprehensive income – Translation Adjustment

6540

(to record translation adjustment applicable to the differential)

Differential = 167400-(700000*0.18) = 41400

Differential allocated of plant property and equipment = $18000

Differential allocated to patent = 41400-18000 = $23400

Differential allocated of plant property and equipment in NKr = 18000/0.18 = $100000

Amortization of plant property and equipment in NKr = 100000/10 = NKr 10000

Amortization of plant property and equipment in $ = 10000*0.20 = $2000

Differential allocated of plant property and equipment in NKr = 23400/0.18 = $130000

Amortization of plant property and equipment in NKr = 130000/5 = NKr 26000

Amortization of plant property and equipment in $ = 26000*0.20 = $5200

(100000*0.18)-(10000*0.20)-((100000-10000)*0.21) = 2900

(130000*0.18)-(26000*0.20)-((130000-26000)*0.21) = 3640

Part C

Income from Pirate’s operations for 20X5, exclusive of income from the Norwegian subsidiary

254000

Add: Income from the Norwegian subsidiary for 20X5 (791000-411000-120000-67000)*0.20

38600

Deduct: Amortization of differential

(7200)

Pirate’s Net Income

285400

Add: Translation Adjustment ($22110+6540)

28950

Pirate’s Consolidated Comprehensive Income

$314350

Part D

Pirate’s stockholders’ equity at Jan. 1, 20X5

3600000

Add: Net income for 20X5

314350

Deduct: Dividends declared by Pirate during 20X5

(130000)

Add: Accumulated other comprehensive income:

Foreign currency translation adjustment

28950

Consolidated Stockholders’ Equity at Dec. 31, 20X5

$3813300

[P1]